an urn contains 4 red, 2 white, and 4 blue marbles. In how many ways can three marbles be selected so that at least one of them is blue?
Answer by stanbon(75887)

an urn contains 4 red, 2 white, and 4 blue marbles. In how many ways can three marbles be selected so that at least one of them is blue?
-------------------------------
P(no blue) + P(at least one blue) = 1
---------------------------------------------
So, P(at least one blue) = 1 - P(no blue)
---
P(none blue) = 6C3/10C3 = 20/120 = 1/6
----------------
Therefore: P(at least one blue) = 1-(1/6) = 5/6
==================================================
Cheers,
Stan H.